# Narrow linewidth exciton-polariton laser

### v1

@@ Abstract
We present a high-resolution spectroscopic investigation of the energy and linewidth of an exciton-polariton laser in the single-mode regime, which derives its coherent emission from an optically pumped and confined exciton-polariton condensate. We report an ultra-narrow linewidth of 56 MHz or 0.24 $\mu$eV, corresponding to a coherence time of 5.7 ns. The narrow linewidth is consistently achieved by using an exciton-polariton condensate with a high photonic content confined in an optically induced trap. Contrary to previous studies, we show that the excitonic reservoir created by the pump and responsible for creating the trap does not strongly affect the emission linewidth as long as the condensate is trapped and the pump power is well above the condensation (lasing) threshold. Our results open up opportunities for manipulating the macroscopic quantum state of the system owing to its long coherence time.
@@ 1 Introduction
Polariton lasers are coherent light sources generated by the decay of bosonic condensates of exciton-polaritons (polaritons for short)  [3, 4], hybrid particles arising from the strong coupling of photons and excitons (electron-hole pairs). These lasers, typically built using an optical microcavity with embedded exciton-hosting material, achieve lasing operation without population inversion [5] and at a threshold much lower than that of an equivalent conventional photon laser [6]. Their low-threeshold operation, hybrid light-matter nature, large nonlinearity, and solid-state platform are promising for a number of low-power applications, such as ultra-fast optical polarization switches, modulation applications, compact sources of terahertz radiation, and logic elements [7, 8, 9]. Polariton condensates also find applications in both classical and quantum computing [10] where a long coherence time and a narrow linewidth of the polariton emission can play an important role.
Figure 1: Experimental configuration. (a) Schematics of the experiment showing the excitation laser profile with parameters height $H$, inner diameter $D$ and inner offset $H_{0}$, which results in a similarly-shaped optically-induced trap $\mathcal{V}$ for the polariton condensate. (b) Example real space condensate emission (bright emission) inside the trap created by the pump laser (dashed line corresponds to the edges of the pump). (c) Momentum-space emission of the single-mode condensate. (d) Example spectrum (dots) measured using a scanning Fabry-Pérot interferometer (SFPI) fitted to a Lorentzian function with a FWHM of $\ell_{\nu}=56$ MHz.
@@ 2 Excitation profile and linewidth measurement
We use a ring-shaped excitation profile to create and trap the condensate [26, 27, 28], which has been an established trapping configuration for studying the fundamental properties of polariton condensates, such as the temporal coherence [18], thermalization [29], excitation spectra [30, 31, 32, 33], and critical fluctuations [34]. As schematically drawn in Fig. 1(a), the excitation profile has an annular shape, with inner diameter $D$, wall thickness $W$, wall height $H$, and inner offset $H_{0}$. The resulting optically induced potential $\mathcal{V}$ assumes a similar shape but with ‘soft’ edges due to the excitonic reservoir diffusion  [26].
@@
Figure 2: Pump-power dependence. (a) Intensity and (b) energy of the condensate as a function of normalized excitation power, $P/P_{th}$. The slopes of the linear fits in (b) are 0.0185 meV(or 4.35 GHz)$/(P/P_{th})$ and 0.111 meV(or 26.8 GHz)$/(P/P_{th})$ for the black and red lines, respectively. (c) Linewidth of the polariton emission as a function of $P/P_{th}$ measured using (black dots) the spectrometer and (blue box plots) the SFPI. The box plots show the interquartile range (length of box), the median (red line), and the whole range (dashed lines). The dots are considered outliers (see SI for details). Left inset: Example measured signal from the SFPI of a single quasi-CW condensate pulse scanned over 3 FSRs. Right inset: Combined measurements in log scale. (d-f) Distribution of the measured linewidths at $P=$1.04 $P_{th}$, 1.16 $P_{th}$, and 1.80 $P_{th}$, respectively. The standard deviations are (d) 150 MHz, (e) 40 MHz, and (f) 8 MHz. (g) Standard deviation of linewidth, $\Delta\ell_{v}$. (h) Energy distributions at $P=1.16P/P_{th}$ measured (red) between shots, centered with respect to the median, i.e. $E-E_{med}$, and (green) within the shot between two FSRs, i.e. $E-E_{FSR}$. The standard deviations are (red) $0.8\leavevmode\nobreak\ GHz$ and (green) $0.64$ GHz. (i) Scattered plot of simultaneous linewidth and energy measurements corresponding to (f).
The sample is a high Q-factor GaAs/AlGaAs planar microcavity with a cavity photon lifetime, $\tau_{C}\sim$150 ps [35], enclosed in a cryostat at a temperature of $<$10 K. We restrict the measurements to a region of the sample where polaritons have low excitonic (or high photonic) content at an excitonic fraction of $|X|^{2}$=$0.13$, corresponding to a detuning of $\delta$=$-17$ meV and a polariton lifetime of $\sim 175\leavevmode\nobreak\ ps$ (see SI for more details). We use the low excitonic fraction to minimise strong linewidth broadening due to interactions [12, 2, 17, 13] while allowing for the formation of a single-mode confined condensate in the ground state. If the excitonic fraction is too low, the condensate tends to occupy several high-energy states of the trap [27, 36, 37]. An example of polariton condensate photoluminescence (PL) emission is presented in Fig. 1(b) showing the confined condensate in the center of the ring-shape potential barrier created by the excitation laser. The corresponding energy and momentum distribution is presented in Fig. 1(c) showing a single-mode ground-state emission.
We use a SFPI from Thorlabs (SA210-5B) with a 10-GHz free spectral range (FSR) to measure the spectrum of the single-mode polariton laser (see SI for details). We synchronized the chopping of our laser and the scanning of the Fabry-Pérot such that at least 3 FSRs are spanned in every single quasi-CW pulse of the condensate. An example of the measured spectrum fitted to a Lorentzian function is presented in Fig. 1(d).
@@ 3 Excitation power dependence
where $E_{0}$ is the polariton energy at zero momentum in the low-density limit. The second term $\mathcal{T}_{0}(P)$ is the quantum confinement energy of the optically induced trap, which increases with $P$ since the trap gets tighter and potential walls higher with increasing $P$ [39]. The last two terms originate from interparticle interactions, where $g$ ($g_{R}$) is the polariton–polariton (polariton–reservoir) interaction strength and $n_{c,0}$ ($n_{R,0}$) is the condensate (reservoir) density around the center of the trap. Both the reservoir and condensate densities grows with $P$, such that at high excitation powers, $P>P_{th}$, they dominate over $\mathcal{T}_{0}$. Note that the trapping potential is created by the reservoir distribution, i.e. $\mathcal{V}(x,y)=g_{R}n_{R}(x,y)$. However, we simplify it here such that $\mathcal{V}(x,y)=\mathcal{T}_{0}+\mathcal{V}_{0}$, where $\mathcal{T}_{0}$ captures the shape of the potential barrier and $\mathcal{V}_{0}=g_{R}n_{R,0}$ captures the energy shifts imposed by the trap offset $\mathcal{V}_{0}$ due to non-negligible reservoir density at the bottom of the trap. Note that $\mathcal{V}_{0}$ is proportional to the inner offset $H_{0}$ of the excitation profile, while $\mathcal{T}_{0}$ is strongly affected by the trap diameter $D$. Equation (3) clearly suggests that any fluctuation in both the reservoir $n_{R}$ and condensate $n_{c}$ densities leads to energy fluctuations of the condensate, and hence to the linewidth broadening.
@@
The resulting distributions of measured linewidths are presented in Fig. 2(c) using a box plot to simultaneously show the asymmetry and spread of the data. The plot clearly shows that the linewidth continues to narrow as the excitation power is increased. For $P/P_{th}>1.2$, the linewidths plateau to a median of $\approx$100 MHz up to the maximum excitation power of $P_{max}/P_{th}\approx 1.8$. This trend is accompanied by a narrowing of the spread of the shot-to-shot variations of the linewidth, down to a standard deviation of $\Delta\ell_{\nu}=8\leavevmode\nobreak\ MHz$, as shown in Fig. 2(g). The latter suggests that the linewidth is becoming less sensitive to fluctuations at higher excitation powers. As shown in the next section, the same trend persists as we change other parameters of the trap that have a direct effect on the overlap of the condensate and the reservoir. Note that close to the maximum power of $P/P_{th}=1.8$, the distribution of linewidths are relatively symmetric [see Fig. 2(f)] and the fluctuations in linewidth is not correlated to the fluctuations in energy [see Fig. 2(i)].
@@ 4 Dependence on trap parameters
We first tune the inner offset $H_{0}$ while keeping all other parameters of the excitation profile fixed (see SI for more details). This offset is proportional to the third term ($\mathcal{V}_{0}=g_{R,0}n_{R}$) in Eq. (3) and will not strongly affect the other terms. Hence, this measurement directly probes the influence of the reservoir on the polariton laser linewidth. The dependence of the linewidth on $H_{0}$ at the maximum excitation laser power is presented in Fig. 3(a). Interestingly, the offset does not significantly affect the linewidth ($\sim$ 100 MHz) until a certain value, at which the linewidth drastically broadens to $\sim$ 1 GHz. Note that increasing $H_{0}$ also increases the excitation power $P$. However, the observed dependence on $H_{0}$ is the reverse of the $P/P_{th}$ dependence observed in Fig. 2(c), where the linewidth drastically decreases with $P$. This is because changing $H_{0}$ unavoidably changes $P_{th}$.
@@
To compare the results on an equal footing, we scanned $P$ for different offsets $H_{0}$ to determine $P_{th}$ and the results are presented in Fig. 3(b) using the normalized excitation power $P/P_{th}$. The trend clearly shows a general behavior consistent across different offsets: the linewidth narrows with increasing $P/P_{th}$ and plateaus to $\sim$100 MHz at $P/P_{th}>1.2$. It now matches the power dependence in Fig. 2(c). However, larger offsets tend to have larger linewidths at the same normalized excitation power, especially for $P/P_{th}<1.2$. In addition, the linewidth variations, $\Delta\ell_{\nu}$, shown in Fig. 3(c), decrease and plateau in a similar manner but at $P/P_{th}>1.5$. These findings strongly suggest that the overlap between the excitonic reservoir and condensate only plays a significant role in linewidth broadening near the condensation threshold, i.e. for $P<1.2P_{th}$. At higher excitation powers, the trend is the same regardless of the trap offset (at least within our spectral resolution), and the linewidth is robust against power fluctuations. This could be due to the hole-burning effect [42, 27] that results in lower than expected trap offsets $\mathcal{V}_{0}$ at higher excitation powers, regardless of the excitation offset $H_{0}$.
@@
Note that tuning $H_{0}/H$ towards large values also allows us to probe the interplay between trapping and localization by gain [13], since the trapping becomes less effective for large offsets ($H_{0}/H>0.45$) such that an “untrapped” condensate [20] can form at $H_{0}/H\rightarrow 0$. In previous experiments [18], untrapped condensates tends to have order-of-magnitude larger linewidths (short coherence times) than trapped ones. Hence, the significantly large linewidths observed for large offsets ($H_{0}/H>0.45$) can be also be attributed to the increased role of gain [13].
@@
We also tune the inner diameter $D$ of the trap while fixing the excitation intensity, the wall width $W$, and inner offset ($H_{0}=0$). Note, however, that the parameter $V$, which is proportional to the potential height, changes with $D$ since the amplitude mask will sample different radial distances from the center of the input Gaussian beam [see Fig. 3(d) inset]. Changing $D$ will mainly affect $\mathcal{T}_{0}$ in Eq. (3), since the confinement energy is inversely proportional to the trap area. The change will also affect $n_{R,0}$ but this effect will only be strong when $D$ is small, i.e. when the center of the trap is closer to the walls. In both cases, we expect the energy fluctuations to be stronger, and hence linewidths broader, for smaller $D$.
@@
The non-monotonic behavior arises from the unavoidable change of the excitation power $P$ and condensation threshold $P_{th}$ as $D$ is varied. The normalized excitation power $P/P_{th}$, plotted as orange dots in Fig. 3(d), clearly shows the same (albeit inverted) non-monotonic behavior. This provides further evidence that the linewidth tends to broaden close to threshold. However, for the same $P/P_{th}$, smaller traps tend to have a larger median linewidths [see Fig. 3(f)]. For example, for the similar ratio of $P/P_{th}\approx 1.1$, the condensate in the 6-$\mu$m trap has a median $\ell_{\nu}\approx 1$ Ghz, which is an order of magnitude larger than that in the 16-${\mu m}$ trap. This behavior is similar to Fig. 3(b), further suggesting that the overlap with the reservoir, which is large for smaller traps, strongly broadens the linewidth at excitation powers close to threshold.
In addition to changing $P_{th}$, increasing the trap size can promote multi-mode behavior due to increased occupation of the excited states as the level spacing decreases. Spontaneous emission into these states effectively results in additional noise to the lasing mode (the ground state) [13], which can contribute the observed broadening for $D>12\mu$m.

### v2

# Narrow-linewidth exciton-polariton laser

@@ Abstract
Exciton-polariton laser is a promising source of coherent light for low-energy applications due to its low-threshold operation. However, a detailed experimental study of its spectral purity, which directly affects its coherence properties is still missing. Here, we present a high-resolution spectroscopic investigation of the energy and linewidth of an exciton-polariton laser in the single-mode regime, which derives its coherent emission from an optically pumped and confined exciton-polariton condensate. We report an ultra-narrow linewidth of 56 MHz or 0.24 $\mu$eV, corresponding to a coherence time of 5.7 ns. The narrow linewidth is consistently achieved by using an exciton-polariton condensate with a high photonic content confined in an optically induced trap. Contrary to previous studies, we show that the excitonic reservoir created by the pump and responsible for creating the trap does not strongly affect the emission linewidth as long as the condensate is trapped and the pump power is well above the condensation (lasing) threshold. The long coherence time of the exciton-polariton system uncovered here opens up opportunities for manipulating its macroscopic quantum state, which is essential for applications in classical and quantum computing.
@@ 1 Introduction
Polariton lasers are coherent light sources generated by the decay of bosonic condensates of exciton-polaritons (polaritons for short)  [3, 4], hybrid particles arising from the strong coupling of photons and excitons (electron-hole pairs). These lasers, typically built using an optical microcavity with embedded exciton-hosting material, achieve lasing operation without population inversion [5] and at a threshold much lower than that of an equivalent conventional photon laser [6]. Their low-threshold operation, hybrid light-matter nature, large nonlinearity, and solid-state platform are promising for a number of low-power applications, such as ultra-fast optical polarization switches, modulation applications, compact sources of terahertz radiation, and logic elements [7, 8, 9]. Polariton condensates also find applications in both classical and quantum computing [10] where a long coherence time and a narrow linewidth of the polariton emission can play an important role.
Figure 1: Experimental configuration. (a) Schematics of the experiment showing the excitation laser profile with parameters height $H$, inner diameter $D$ and inner offset $H_{0}$, which results in a similarly-shaped optically-induced trap $\mathcal{V}$ for the polariton condensate. (b) Example real space condensate emission (bright emission) inside the trap created by the pump laser (dashed line corresponds to the edges of the pump). (c) Momentum-space emission of the single-mode condensate. The dashed line is the polariton dispersion extracted from low-power measurement (see Supplemental Document). (d) Example spectrum (black dots) of the emission of the condensate or the polariton laser measured using the scanning Fabry-Pérot Interferometer (SFPI) and fitted to a Lorentzian function (red solid line) with a FWHM of $\ell_{\nu}=56$ MHz. There is no other peak within the free spectral range of the SFPI (see Supplemental Document). Here, the excitation and condensate emission powers are $48.6$ mW and 60 $\mu$W, respectively.
@@ 2 Excitation profile and linewidth measurement
We use a ring-shaped excitation profile to create and trap the condensate [26, 27, 28], which has been an established trapping configuration for studying the fundamental properties of polariton condensates, such as the temporal coherence [18], thermalization [29, 30], excitation spectra [31, 32, 33, 34], and critical fluctuations [35]. As schematically drawn in Fig. 1(a), the excitation profile has an annular shape, with inner diameter $D$, wall thickness $W$, wall height $H$, and inner offset $H_{0}$. The resulting optically induced potential $\mathcal{V}$ assumes a similar shape but with ‘soft’ edges due to the excitonic reservoir diffusion  [26].
@@
Figure 2: Pump-power dependence. (a) Intensity and (b) energy of the condensate as a function of normalized excitation power, $P/P_{th}$. The slopes of the linear fits in (b) are 0.0185 meV(or 4.35 GHz)$/(P/P_{th})$ and 0.111 meV(or 26.8 GHz)$/(P/P_{th})$ for the black and red lines, respectively. (c) Linewidth of the polariton emission as a function of $P/P_{th}$ measured using (black dots) the spectrometer and (blue box plots) the SFPI. The box plots show the interquartile range (length of box), the median (red line), and the whole range (dashed lines). The dots are considered outliers (see SI for details). Left inset: Example measured signal from the SFPI of a single quasi-CW condensate pulse scanned over 3 FSRs of the SFPI. Right inset: Combined measurements in log scale. (d-f) Distribution of the measured linewidths at $P=$1.04 $P_{th}$, 1.16 $P_{th}$, and 1.80 $P_{th}$, respectively. The standard deviations are (d) 150 MHz, (e) 40 MHz, and (f) 8 MHz. (g) Standard deviation of linewidth, $\Delta\ell_{v}$. (h) Energy distributions at $P=1.16P/P_{th}$ measured (red) between shots, centered with respect to the median, i.e. $E-E_{med}$, and (green) within the shot between two FSRs, i.e. $E-E_{FSR}$. The standard deviations are (red) $0.8\leavevmode\nobreak\ GHz$ and (green) $0.64$ GHz. (i) Scattered plot of simultaneous linewidth and energy measurements corresponding to (f). For this dataset, $P_{th}\approx 30$ mW.
[…]
We use a SFPI from Thorlabs (SA210-5B) with a 10-GHz free spectral range (FSR) to measure the spectrum of the single-mode polariton laser (see Supplemental Document for details). We synchronized the chopping of our laser and the scanning of the Fabry-Pérot such that at least 3 FSRs are spanned in every single quasi-CW pulse of the condensate. An example of the measured spectrum fitted to a Lorentzian function is presented in Fig. 1(d).
@@ 3 Excitation power dependence
where $E_{0}$ is the polariton energy at zero momentum in the low-density limit. The second term $\mathcal{T}_{0}(P)$ is the quantum confinement energy of the optically induced trap, which increases with $P$ since the trap gets tighter and potential walls higher with increasing $P$ [40]. The last two terms originate from interparticle interactions, where $g$ ($g_{R}$) is the polariton–polariton (polariton–reservoir) interaction strength and $n_{c,0}$ ($n_{R,0}$) is the condensate (reservoir) density around the center of the trap. Both the reservoir and condensate densities grows with $P$, such that at high excitation powers, $P>P_{th}$, they dominate over $\mathcal{T}_{0}$. Note that the trapping potential is created by the reservoir distribution, i.e. $\mathcal{V}(x,y)=g_{R}n_{R}(x,y)$. However, we simplify it here such that $\mathcal{V}(x,y)=\mathcal{T}_{0}+\mathcal{V}_{0}$, where $\mathcal{T}_{0}$ captures the shape of the potential barrier and $\mathcal{V}_{0}=g_{R}n_{R,0}$ captures the energy shifts imposed by the trap offset $\mathcal{V}_{0}$ due to non-negligible reservoir density at the bottom of the trap. Note that $\mathcal{V}_{0}$ is proportional to the inner offset $H_{0}$ of the excitation profile, while $\mathcal{T}_{0}$ is strongly affected by the trap diameter $D$. Equation (3) clearly suggests that any fluctuation in both the reservoir $n_{R}$ and condensate $n_{c}$ densities leads to energy fluctuation of the condensate, and hence to the linewidth broadening.
@@
The resulting distributions of measured linewidths are presented in Fig. 2(c) using a box plot to simultaneously show the asymmetry and spread of the data. The plot clearly shows that the linewidth continues to narrow as the excitation power is increased. For $P/P_{th}>1.2$, the linewidths plateau to a median of $\approx$100 MHz up to the maximum excitation power of $P_{max}/P_{th}\approx 1.8$. This trend is accompanied by a narrowing of the spread of the shot-to-shot variations of the linewidth, down to a standard deviation of $\Delta\ell_{\nu}=8\leavevmode\nobreak\ MHz$, as shown in Fig. 2(g). The latter suggests that the linewidth is becoming less sensitive to fluctuations at higher excitation powers. As shown in the next section, the same trend persists as we change other parameters of the trap that have a direct effect on the overlap of the condensate and the reservoir. Note that close to the maximum power of $P/P_{th}=1.8$, the distribution of linewidths are relatively symmetric [see Fig. 2(f)] and the fluctuation in linewidth is not correlated to the fluctuation in energy [see Fig. 2(i)].
@@ 4 Dependence on trap parameters
We first tune the inner offset $H_{0}$ while keeping all other parameters of the excitation profile fixed (see Supplemental Documen for more details). This offset is proportional to the third term ($\mathcal{V}_{0}=g_{R,0}n_{R}$) in Eq. (3) and will not strongly affect the other terms. Hence, this measurement directly probes the influence of the reservoir on the polariton laser linewidth. The dependence of the linewidth on $H_{0}$ at the maximum excitation laser power is presented in Fig. 3(a). Interestingly, the offset does not significantly affect the linewidth ($\sim$ 100 MHz) until a certain value, at which the linewidth drastically broadens to $\sim$ 1 GHz. Note that increasing $H_{0}$ also increases the excitation power $P$. However, the observed dependence on $H_{0}$ is the reverse of the $P/P_{th}$ dependence observed in Fig. 2(c), where the linewidth drastically decreases with $P$. This is because changing $H_{0}$ unavoidably changes $P_{th}$.
@@
To compare the results on an equal footing, we scanned $P$ for different offsets $H_{0}$ to determine $P_{th}$ and the results are presented in Fig. 3(b) using the normalized excitation power $P/P_{th}$. The trend clearly shows a general behavior consistent across different offsets: the linewidth narrows with increasing $P/P_{th}$ and plateaus to $\sim$100 MHz at $P/P_{th}>1.2$. It now matches the power dependence in Fig. 2(c). However, larger offsets tend to have larger linewidths at the same normalized excitation power, especially for $P/P_{th}<1.2$. In addition, the linewidth variations, $\Delta\ell_{\nu}$, shown in Fig. 3(c), decrease and plateau in a similar manner but at $P/P_{th}>1.5$. These findings strongly suggest that the overlap between the excitonic reservoir and condensate only plays a significant role in linewidth broadening near the condensation threshold, i.e. for $P<1.2P_{th}$. At higher excitation powers, the trend is the same regardless of the trap offset (at least within our spectral resolution), and the linewidth is robust against power fluctuations. This could be due to the hole-burning effect [43, 27] that results in lower than expected trap offsets ${H}_{0}$ at higher excitation powers, regardless of the excitation offset $H_{0}$.
@@
zzzNote that tuning $H_{0}/H$ towards large values also allows us to probe the interplay between trapping and localization by gain [13], since the trapping becomes less effective for large offsets ($H_{0}/H>0.45$) such that an “untrapped” condensate [20] can form at $H_{0}/H\rightarrow 1$. In previous experiments [18], untrapped condensates tends to have order-of-magnitude larger linewidths (short coherence times) than trapped ones. Hence, the significantly large linewidths observed for large offsets ($H_{0}/H>0.45$) can be also be attributed to the increased role of gain [13].
@@
We also tune the inner diameter $D$ of the trap while fixing the excitation intensity, the wall width $W$, and inner offset ($H_{0}=0$). Note, however, that the parameter $H$, which is proportional to the potential height, changes with $D$ since the amplitude mask will sample different radial distances from the center of the input Gaussian beam [see Fig. 3(d) inset]. Changing $D$ will mainly affect $\mathcal{T}_{0}$ in Eq. (3), since the confinement energy is inversely proportional to the trap area. The change will also affect $n_{R,0}$ but this effect will only be strong when $D$ is small, i.e. when the center of the trap is closer to the walls. In both cases, we expect the energy fluctuation to be stronger, and hence linewidths broader, for smaller $D$.
@@
The non-monotonic behavior arises from the unavoidable change of the excitation power $P$ and condensation threshold $P_{th}$ as $D$ is varied. The normalized excitation power $P/P_{th}$, plotted as orange dots in Fig. 3(d), clearly shows the same (albeit inverted) non-monotonic behavior. This provides further evidence that the linewidth tends to broaden close to threshold. However, for the same $P/P_{th}$, smaller traps tend to have larger median linewidths [see Fig. 3(f)]. For example, for the similar ratio of $P/P_{th}\approx 1.1$, the condensate in the 6-$\mu$m trap has a median $\ell_{\nu}\approx 1$ Ghz, which is an order of magnitude larger than that in the 16-${\mu m}$ trap. This behavior is similar to Fig. 3(b), further suggesting that the overlap with the reservoir, which is large for smaller traps, strongly broadens the linewidth at excitation powers close to threshold.
In addition to changing $P_{th}$, increasing the trap size can promote multi-mode behavior due to increased occupation of the excited states as the level spacing decreases. Spontaneous emission into these states effectively results in additional noise to the lasing mode (the ground state) [13], which can contribute to the observed broadening for $D>12\mu$m.